# Trace anomaly contributions to semi-classical wormhole geometries

### v1

@@ II.2 Morris-Thorne wormhole geometries: Semi-classical approach
as a possible solution to obtain viable traversable wormhole geometries. In the metric presented above, $\Phi(r)$ denotes the redshift function, which determines the gravitational redshift experienced by signals traversing the spacetime. The function $b(r)$ is referred to as the wormhole shape function, as it characterizes the spatial geometry of the wormhole and determines its spatial embedding. For the wormhole to possess a physically viable geometry, the shape function must satisfy the flaring-out condition at the throat. Specifically, at the throat radius $r_{0}$, one must have $b^{\prime}(r_{0})<1$, and throughout the entire spacetime region $r>r_{0}$, the condition $b(r)<r$ must hold. These requirements ensure that the wormhole throat is properly sustained and that the spacetime remains free from horizons or singularities in the traversable region.
@@ III.1 Specific case: constant redshift function
$g(r)=\frac{-4r^{2}+8\lambda}{8\lambda}+\frac{r^{2}}{8\lambda}\,{\left(2W\left[\pm\frac{\sqrt{2C_{1}\lambda}}{r^{3}}+2\right]\right)}^{2},$ (22)
where $W[x]$ denotes the principal solution of the Lambert function [81, 82]. For real arguments, the Lambert function can be viewed as the inverse of the relation $x=W(x)\,e^{W(x)}$ [83].
@@
which guarantees the asymptotic flatness of the solution. Considering different values of the parameter $\lambda$, we plot the function 1b(r)/r1-b(r)/r in Fig. (1).
@@
In order to analyze the energy conditions for this class of solutions, we study the behavior of the quantities $\rho$, $\rho+p_{r}$, and $\rho+p_{t}$ throughout the spacetime, which are expressed as:
$\rho=-\frac{3}{2\lambda}\,{W\left[-\frac{r_{0}^{2}{[-2\lambda+2r_{0}\zeta]}^{1/2}}{\lambda r^{3}}\right]}^{2},$ (27)
$\rho+p_{r}=-\frac{1}{\lambda}\,{W\left[-\frac{r_{0}^{2}{[-2\lambda+2r_{0}\zeta]}^{1/2}}{\lambda r^{3}}\right]}\left({W\left[-\frac{r_{0}^{2}{[-2\lambda+2r_{0}\zeta]}^{1/2}}{\lambda r^{3}}\right]}-1\right),$ (28)
$\rho+p_{t}=-\frac{\left({W\left[-\frac{r_{0}^{2}{[-2\lambda+2r_{0}\zeta]}^{1/2}}{\lambda r^{3}}\right]}\right)\left(3{W\left[-\frac{r_{0}^{2}{[-2\lambda+2r_{0}\zeta]}^{1/2}}{\lambda r^{3}}\right]}+2{W\left[-\frac{r_{0}^{2}{[-2\lambda+2r_{0}\zeta]}^{1/2}}{\lambda r^{3}}\right]}^{2}+1\right)}{2\lambda\left({W\left[-\frac{r_{0}^{2}{[-2\lambda+2r_{0}\zeta]}^{1/2}}{\lambda r^{3}}\right]}+1\right)},$ (29)
respectively, and in the asymptotic limit, reduce to:
@@ III.2.1 Curvature coordinates
$-4\,{r}^{4}\lambda\,{f}^{2}\left(r-{r_{0}}\right)^{2}{{f^{\prime\prime}}}^{2}-{r}^{4}\lambda\,\left(r-{r_{0}}\right)^{2}{{f^{\prime}}}^{4}-\left(-12{r_{0}}+4r\right)f{r}^{3}\left(r-{r_{0}}\right)\lambda\,{{f^{\prime}}}^{3}-f^{\prime}f^{2}\,\left(9\lambda-18\alpha\right){{r_{0}}}^{2}$
@@ III.2.2 Isotropic coordinates
$ds^{2}=-\left[h_{0}+h_{1}\left(\frac{1-\frac{r_{0}}{4\tilde{r}}}{1+\frac{r_{0}}{4\tilde{r}}}\right)-\alpha\left(\frac{64\left(1-\frac{r_{0}}{4\tilde{r}}\right)\Xi(r)}{35r_{0}^{2}\left(1+\frac{r_{0}}{4\tilde{r}}\right)^{7}}\right)\right]^{2}dt^{2}+{\left(1+\frac{r_{0}}{4\tilde{r}}\right)}^{4}\left[d{\tilde{r}}^{2}+{\tilde{r}}^{2}\left(d\theta^{2}+\sin^{2}\theta\,d\phi^{2}\right)\right]\,$, (48)
@@
$\Xi(r)=\frac{r_{0}^{6}+32\,\tilde{r}r_{0}^{5}+464\,{\tilde{r}}^{2}r_{0}^{4}+4096\,{\tilde{r}}^{3}{r_{0}}^{3}+7424\,{\tilde{r}}^{4}{{r_{0}}}^{2}+8192\,{\tilde{r}}^{5}{r_{0}}+4096\,{\tilde{r}}^{6}}{4096{\tilde{r}}^{6}}.$ (49)
@@
One verifies that the area attains its minimum at $\tilde{r}=r_{0}/4$, and the flaring-out condition is satisfied, namely $A^{\prime\prime}\!\left(\tilde{r}=\frac{r_{0}}{4}\right)=64\pi>0$. Therefore, the geometry corresponds to a traversable wormhole with the throat located at $\tilde{r}=r_{0}/4$, in the absence of an event horizon. An event horizon would arise if $g_{tt}$ vanishes, leading to the horizon solution given by
@@
We verify that the Ricci scalar (17) and the Kretschmann scalar (18) diverge at $\tilde{r}_{H}$ for the global coordinate patch given in Eq. (48), indicating the presence of a naked singularity at $\tilde{r}_{H}$. However, a curvature singularity does not form in the physical region, since the $g_{tt}$ component of the metric never vanishes. Consequently, the solution describes a traversable wormhole in the regions where $\tilde{r}_{H}$ is negative. In the limit $\alpha\to 0$, we obtain
@@ III.3 General case: $\lambda$ and $\alpha$
In this section, we solve the differential equation (36) to obtain an exact solution at or near the wormhole throat. This approach ensures that the redshift function remains finite in the vicinity of the traversable wormhole throat. We focus on the radial region $r_{0}<r<r_{0}+\Delta$, where $\Delta$ is taken to be infinitesimally small. In this near-throat region, Eq. (36) can be simplified as follows:
$\lambda r^{2}f^{\prime\prime}(r)-6r\left(r^{2}+2\lambda-4\alpha\right)f(r)f^{\prime}(r)+36\lambda f^{2}(r)=0.$ (58)
We perform a transformation $f(r)={\rm exp}(W(r))$ which yields
$r^{2}\lambda{W^{\prime}(r)}^{2}-6r\left(r^{2}+2\lambda-4\alpha\right)W^{\prime}(r)+36\lambda=0,$ (59)
which can easily be solved for the first derivative of $W(r)$ with respect to $r$
$W^{\prime}(r)=\frac{3\left(r^{2}+2\lambda-4\alpha\right)\pm\sqrt{\left(r^{2}-4\alpha\right)\left(r^{2}+4\lambda-4\alpha\right)}}{\lambda r}.$ (60)
An interesting case arises when the negative sign is chosen in the above solutions, which corresponds to an asymptotically flat spacetime. In this case, we can solve for $W(r)$ as
$W(r)$ $=$ $\frac{1}{\lambda}\left[2\left(2\alpha-\lambda\right)\ln\left(\frac{r^{2}+2\lambda-4\alpha+\gamma_{0}}{\sqrt{r}}\right)+r^{2}-\gamma_{0}\right]$ (61)
$+\frac{4\sqrt{\alpha(\alpha-\lambda)}}{\lambda}\,\ln\left[\frac{8\gamma_{0}\sqrt{\alpha(\alpha-\lambda)}+4r^{2}(\lambda-2\alpha)+32\alpha(\alpha-\lambda)}{r^{2}}\right]+W_{0},$
where
$\gamma_{0}=\sqrt{(r^{2}-4\alpha)(r^{2}+4\lambda-4\alpha)}$ (62)
For $r_{0}^{2}>4\alpha$, we thus find that the Ricci and Kretschmann scalars take the following finite values at the throat:
$\mathcal{R}({r_{0}})=\frac{6\lambda}{r_{0}^{4}}+\frac{24\alpha\lambda-12\lambda^{4}}{r_{0}^{6}}+{\mathcal{O}}\left(\frac{1}{r_{0}^{8}}\right),$ (63)
@@
In Fig. 4, we show the numerical solution of Eq. (36) for the redshift function. It is observed that $f(r)$ remains finite everywhere, as expected, indicating the absence of an event horizon in the wormhole spacetime. Using Eqs. (11) and (12), we then obtain
$p_{r}\simeq\frac{3r^{2}(r-r_{0})-3\gamma_{0}(r-r_{0})-12\alpha(r-r_{0})+\lambda(6r-7r_{0})}{\lambda r^{3}},$ (65)
@@
$p_{t}\simeq\frac{1152\alpha^{3}(r-r_{0})+\alpha^{2}\left[\lambda\left(1680r_{0}-1728r\right)+864r_{0}r^{2}+144\gamma_{0}(r-r_{0})-864r^{3}\right]+\Sigma_{5}}{4\gamma_{0}\lambda^{2}r^{3}},$ (66)
where
$\Sigma_{5}$ $=$ $\lambda^{2}\left(576r-528r_{0}\right)+\lambda\alpha\left(\gamma_{0}(132r_{0}-144r)+912r^{3}-888r_{0}r^{2}\right)+9\gamma_{0}^{3}(r-r_{0})$ (67)
$+\alpha\left[72\gamma_{0}(r_{0}{r}^{2}-r^{3})-216r_{0}{r}^{4}+216r^{5}\right]+\left(156\,{\it r_{0}}\,{r}^{2}-168\,{r}^{3}+\left(36\,r-28\,{\it r_{0}}\right)\gamma_{0}\right){\lambda}^{2}$
$+\left(\left(48\,{r}^{3}-45\,{\it r_{0}}\,{r}^{2}\right)\gamma_{0}-120\,{r}^{5}+117\,{\it r_{0}}\,{r}^{4}\right)\lambda-18\,{r}^{7}+18\,{r}^{6}{\it r_{0}}+\left(9\,{r}^{5}-9\,{\it r_{0}}\,{r}^{4}\right)\gamma_{0}\,.$
@@
In Fig. 5, we illustrate the behavior of the radial and tangential pressures as functions of the radial coordinate for different values of the parameters $\lambda$ and $\alpha$. In particular, the magnitude of the radial pressure increases near the wormhole throat as the parameter $\lambda$ is increased.

### v2

@@ II.2 Morris-Thorne wormhole geometries: Semi-classical approach
as a possible solution to obtain viable traversable wormhole geometries. In the metric presented above, $\Phi(r)$ denotes the redshift function, which determines the gravitational redshift experienced by signals traversing the spacetime. The function $b(r)$ is referred to as the wormhole shape function, as it characterizes the spatial geometry of the wormhole and determines its spatial embedding. The radial coordinate ranges from $r_{0}$ to spatial infinity where the surface at $r=r_{0}$ is known as the wormhole’s throat. Every wormhole configuration must have a minimum radius, i.e. at the throat of the wormhole $b(r_{0})=r_{0}$, where $r_{0}$ being the minimum value of $r$. For the wormhole to possess a physically viable geometry, the shape function must satisfy the flaring-out condition at the throat. Specifically, at the throat radius $r_{0}$, one must have $b^{\prime}(r_{0})<1$, and throughout the entire spacetime region $r>r_{0}$, the condition $b(r)<r$ must hold. These requirements ensure that the wormhole throat is properly sustained and that the spacetime remains free from horizons or singularities in the traversable region.
Note that the wormhole geometry generally requires two coordinate patches, each covering the domain $r\in[r_{0},+\infty[$. Each patch corresponds to one of the two universes, and the patches meet smoothly at $r=r_{0}$, the location of the wormhole throat [4]. While the metric component $g_{rr}=(1-b(r)/r)^{-1}$ diverges as $r\to r_{0}$, this divergence is purely a coordinate effect. One may instead introduce the proper radial distance
$l(r)=\pm\int_{r_{0}}^{r}\frac{dr}{\sqrt{1-b(r)/r}},$ (6)
which remains finite for all finite $r$ throughout the spacetime, so that no thin-shell construction is required at $r_{0}$ (although it may be present [4]). The two signs correspond to the two asymptotically flat regions of the manifold: as $l\to\pm\infty$ (equivalently $r\to\infty$), the ratio $b(r)/r\to 0$, demonstrating that the geometry approaches asymptotic flatness on both sides of the throat.
@@ III.1 Specific case: constant redshift function
$g(r)=\frac{-4r^{2}+8\lambda}{8\lambda}+\frac{r^{2}}{8\lambda}\,{\left(2W\left[\pm\frac{\sqrt{2C_{1}\lambda}}{r^{3}}\right]+2\right)}^{2},$ (23)
where $W[x]$ denotes the principal solution of the Lambert function that is analytic in $x=0$ [81, 82]. For real arguments, the Lambert function can be viewed as the inverse of the relation $x=W(x)\,e^{W(x)}$ [83]. Here, we define $W(x)=W_{0}(x)$ where $W_{0}(x)$ is referred to as the principal branch of the Lambert function. The Taylor series of $W_{0}(x)$ for $|x|<e^{-1}$ is [81]
$W_{0}(x)=\sum_{n=1}^{\infty}{(-n)^{n-1}x^{n}\over n!}.$ (24)
@@
which guarantees the asymptotic flatness of the solution. Considering different values of the parameter $\lambda$, we plot the function 1b(r)/r1-b(r)/r in Fig. (1). To ensure the asymptotic flatness condition, we obtain $l(r)$ by using the solution (28) for small $\lambda$ as
$l(r)=\pm\sqrt{1-\frac{r_{0}}{r}}\bigg(\frac{6r^{2}r_{0}-2\lambda r_{0}-5\lambda r}{6rr_{0}}\bigg)\pm\bigg(\frac{\lambda-2r_{0}^{2}}{4r_{0}}\bigg)\ln\left[\frac{2r}{r_{0}}\left(\sqrt{1-\frac{r_{0}}{r}}+1\right)-1\right],$ (29)
@@
In order to analyze the energy conditions for this class of solutions, we study the behavior of the quantities $\rho$, $\rho+p_{r}$, and $\rho+p_{t}$ throughout the spacetime, which are expressed as
$\rho=-\frac{3}{2\lambda}\,{W\left[-\frac{r_{0}^{2}e^{-\frac{\zeta}{r_{0}}}{[-2\lambda+2r_{0}\zeta]}^{1/2}}{r^{3}}\right]}^{2},$ (30)
$\rho+p_{r}=-\frac{1}{\lambda}\,{W\left[-\frac{r_{0}^{2}e^{-\frac{\zeta}{r_{0}}}{[-2\lambda+2r_{0}\zeta]}^{1/2}}{r^{3}}\right]}\left({W\left[-\frac{r_{0}^{2}e^{-\frac{\zeta}{r_{0}}}{[-2\lambda+2r_{0}\zeta]}^{1/2}}{r^{3}}\right]}-1\right),$ (31)
$\rho+p_{t}=-\frac{\left({W\left[-\frac{r_{0}^{2}e^{-\frac{\zeta}{r_{0}}}{[-2\lambda+2r_{0}\zeta]}^{1/2}}{r^{3}}\right]}\right)\left(3{W\left[-\frac{r_{0}^{2}e^{-\frac{\zeta}{r_{0}}}{[-2\lambda+2r_{0}\zeta]}^{1/2}}{r^{3}}\right]}+2{W\left[-\frac{r_{0}^{2}e^{-\frac{\zeta}{r_{0}}}{[-2\lambda+2r_{0}\zeta]}^{1/2}}{r^{3}}\right]}^{2}+1\right)}{2\lambda\left({W\left[-\frac{r_{0}^{2}e^{-\frac{\zeta}{r_{0}}}{[-2\lambda+2r_{0}\zeta]}^{1/2}}{r^{3}}\right]}+1\right)},$ (32)
respectively, and in the asymptotic limit, these quantities reduce to
@@ III.2.1 Curvature coordinates
$-4\,{r}^{4}\lambda\,{f}^{2}\left(r-{r_{0}}\right)^{2}{{f^{\prime\prime}}}^{2}-{r}^{4}\lambda\,\left(r-{r_{0}}\right)^{2}{{f^{\prime}}}^{4}-\left(-12{r_{0}}+4r\right)f{r}^{3}\left(r-{r_{0}}\right)\lambda\,{{f^{\prime}}}^{3}+4f^{\prime}f^{3}\,\left(9\lambda-18\alpha\right)rr_{0}^{2}$
@@ III.2.2 Isotropic coordinates
$ds^{2}=-\left[h_{0}+h_{1}\left(\frac{1-\frac{r_{0}}{4\tilde{r}}}{1+\frac{r_{0}}{4\tilde{r}}}\right)-\alpha h_{1}\left(\frac{64\left(1-\frac{r_{0}}{4\tilde{r}}\right)\Xi(\tilde{r})}{35r_{0}^{2}\left(1+\frac{r_{0}}{4\tilde{r}}\right)^{7}}\right)\right]^{2}dt^{2}+{\left(1+\frac{r_{0}}{4\tilde{r}}\right)}^{4}\left[d{\tilde{r}}^{2}+{\tilde{r}}^{2}\left(d\theta^{2}+\sin^{2}\theta\,d\phi^{2}\right)\right]\,$, (51)
@@
$\Xi(\tilde{r})=\frac{r_{0}^{6}+32\,\tilde{r}r_{0}^{5}+464\,{\tilde{r}}^{2}r_{0}^{4}+4096\,{\tilde{r}}^{3}{r_{0}}^{3}+7424\,{\tilde{r}}^{4}{{r_{0}}}^{2}+8192\,{\tilde{r}}^{5}{r_{0}}+4096\,{\tilde{r}}^{6}}{4096{\tilde{r}}^{6}}.$ (52)
@@
[…]
@@
To better understand the properties of curvature invariants in these geometries, we can obtain the Ricci scalar (18) and the Kretschmann scalar (19) around the region $\tilde{r}_{H}$ where the $g_{tt}(\tilde{r}_{H})=0$ for the global coordinate patch given in Eq. (51). We therefore find the following expressions for Ricci and Kretschmann scalars in the limit $\tilde{r}\to\tilde{r}_{H}$
$\mathcal{R}(\tilde{r}_{H})=\frac{128{\tilde{r}_{H}}^{4}{g^{\prime 2}_{tt}}(\tilde{r}_{H})}{{g_{tt}(\tilde{r}_{H})}^{2}\left(r_{0}+4\tilde{r}_{H}\right)^{4}},$ (57)
$\mathcal{K}(\tilde{r}_{H})=\frac{16384{\tilde{r}_{H}}^{8}{g^{\prime 4}_{tt}}(\tilde{r}_{H})}{{g_{tt}(\tilde{r}_{H})}^{4}\left(r_{0}+4\tilde{r}_{H}\right)^{8}},$ (58)
whereby we observe that both the curvature scalars diverge as $g_{tt}(\tilde{r}_{H})\to 0$ indicating this horizon is actually a naked curvature singularity. However, the curvature singularity does not form, when the $g_{tt}$ component of the metric (51) never goes to zero, and we therefore have a traversable wormhole. In order to ensure the absence of horizons and singularities, one can choose the parameters characterizing the solution so that $\tilde{r}_{H}<0$. We first examine the limit $\alpha\to 0$ in Eq. (56), and find
@@ III.3 General case: $\lambda$ and $\alpha$
Our objective here is to solve the differential equation (39) for the shape function introduced in the previous section. We then perform a transformation of the form $f(r)={\rm exp}\!\left(2\int W(r)\,dr\right)$ in Eq. (39), which yields
$4\lambda r^{4}\left(r-r_{0}\right)\left[W^{\prime}(r)+W^{2}(r)\right]^{2}-\left(r-r_{0}\right)r^{2}\left(4\lambda\left(2r-3r_{0}\right)rW(r)+6r^{3}-12r_{0}\left(2\alpha-\lambda\right)\right)\left[W^{\prime}(r)+W^{2}(r)\right]$
$+\lambda r^{2}\left(2r-3r_{0}\right)^{2}W^{2}(r)-\big(12r^{4}-9r_{0}r^{3}+12r_{0}r\left(2\alpha-\lambda\right)-18r_{0}^{2}\left(2\alpha-\lambda\right)\big)rW(r)+9\lambda r_{0}^{2}=0.$ (63)
Since Eq. (63) cannot be solved analytically for $W(r)$ in terms of standard functions, we first analyze this equation in the vicinity of the wormhole throat. We restrict our attention to the radial region $r_{0}<r<r_{0}+\Delta$, where $\Delta$ is taken to be infinitesimally small. In this near-throat region, Eq. (63) can be simplified as follows:
$r_{0}^{2}\lambda{W(r_{0})}^{2}-3r_{0}\left(r_{0}^{2}+2\lambda-4\alpha\right)W(r_{0})+9\lambda=0,$ (64)
which can easily be solved for the value of $W(r_{0})$
$W(r_{0})=\frac{3\left(r_{0}^{2}+2\lambda-4\alpha\right)\pm 3\sqrt{\left(r_{0}^{2}+2\lambda-4\alpha\right)^{2}-4\lambda^{2}}}{2\lambda r_{0}}.$ (65)
As the behavior of the redshift function remains finite in the vicinity of the wormhole throat, we expect the Ricci and Kretschmann scalars to attain finite values at the throat. This behavior can be explicitly verified from Eqs. (18) and (19) by considering the negative-sign branch of the above solutions
$\mathcal{R}({r_{0}})=\frac{6\lambda}{r_{0}^{4}}+\frac{24\alpha\lambda-12\lambda^{2}}{r_{0}^{6}}+{\mathcal{O}}\left(\frac{1}{r_{0}^{8}}\right),$ (66)
@@
Also for positive sign solutions the above scalars read
$\mathcal{R}({r_{0}})=\frac{6}{\lambda}+\frac{12\lambda-24\alpha}{\lambda r_{0}^{2}}+{\mathcal{O}}\left(\frac{1}{r_{0}^{4}}\right),$ (68)
@@
$\mathcal{K}({r_{0}})\!\!$ $=$ $\!\!\frac{9}{\lambda^{2}}+\frac{36\lambda-72\alpha}{\lambda^{2}r_{0}^{2}}+{\mathcal{O}}\left(\frac{1}{r_{0}^{4}}\right).$ (69)
It is therefore evident that the Ricci and Kretschmann scalars assume finite values at the throat for the negative-sign branch and remain finite as $\alpha$ and $\lambda$ tend to zero. However, these scalars diverge in the limit $\lambda\to 0$; hence, one may only consider the solutions with the negative sign in Eq. (65). As noted above, obtaining exact analytical solutions of Eq. (63) is extremely challenging. We therefore solve this equation numerically for selected values of the parameters $\alpha$ and $\lambda$ and investigate the corresponding properties of the solutions. In Fig. 4, we display the redshift function obtained using the numerical solution of Eq. (63) for the function $W(r)$, which satisfies the initial condition $W(r_{0})$ at the throat. It is worth noting that $W(r_{0})$ in Eq. (65) with the negative sign yields solutions corresponding to an asymptotically flat spacetime. We observe that $f(r)$ remains finite throughout the spacetime, as expected, indicating the absence of an event horizon in the wormhole geometry. Using Eqs. (12) and (13), we then obtain the following expressions for the radial and transverse pressures:
$p_{r}=\frac{2r\left(r-r_{0}\right)W(r)-r_{0}}{r^{3}},$ (70)
and
$p_{t}=\frac{2r^{2}\left(r-r_{0}\right)W^{\prime}(r)+\big(2r\left(r-r_{0}\right)W(r)+r_{0}\big)\big(1+rW(r)\big)}{2r^{3}}.$ (71)
At wormhole throat, these pressures take the form
$p_{r}\simeq\frac{-1}{r_{0}^{2}}+\bigg(\frac{3+2r_{0}W(r_{0})}{r_{0}^{3}}\bigg)\left(r-r_{0}\right)+{\mathcal{O}}\left(r-r_{0}\right)^{2},$ (72)
and
$p_{t}\simeq\frac{1+r_{0}W(r_{0})}{2r_{0}^{2}}+\bigg(\frac{2r_{0}^{2}W^{2}(r_{0})+3r_{0}^{2}{W^{\prime}}(r_{0})-3}{2r_{0}^{3}}\bigg)\left(r-r_{0}\right)+{\mathcal{O}}\left(r-r_{0}\right)^{2}.$ (73)
In Fig. 5, we illustrate the behavior of the radial and tangential pressures as functions of the radial coordinate by employing Eqs. (70)–(71) together with the numerical solutions for the function $W(r)$, for different values of the parameters $\lambda$ and $\alpha$. In particular, the magnitude of the radial pressure is observed to increase in the vicinity of the wormhole throat as the parameter $\lambda$ increases.